Hello,
I am really struggling with a few of these equations, could you please help me out here and explain how you solved the problem? Thank you very much for your help and assistance with this matter.
Dad of 5!
1. -4+46=7(-3t+6)
2. -4(3b+9)=24+3(2b-8)

Answer by askmemath(368)

1. -4+46=7(-3t+6)
-4 +46 = +42 (Since 46 is the larger number)
7(-3t+6)
In order to remove the brackets,multiply each term within the brackets with the term immediately outside the bracket
Therefore you get -21t + 42
Subtracting 42 on both sides
-21t = 0
which means t = 0
2. -4(3b+9)=24+3(2b-8)
-12b -36 = 24 + 6b -24
-12b -36 = 6b
Adding 12b on both sides
18b = -36
Dividing by 18 on both sides
b = -2